OpenStudy (anonymous):

A vending machine sells chips at $0.40 and candy at $0.50. Last month, the vending machine yielded $152.80 with the sale of 328 items. Let X be the number of chips sold and Y the number of candies. How many packages of chips were purchased last month? @agent0smith

OpenStudy (p0sitr0n):

0.4x+0.5y=152.8 x+y=328 x=328-y .4(328-y)+0.5y=152.8 131.2+.1y=152.8 21.6=.1y 216=y 328-216=112 x=112

OpenStudy (anonymous):

Thanks :)